OpenStudy (anonymous):

when joe is 12 years older, he will be 36. How old is he now? (i suck at word problems and i need help writing an equation)

OpenStudy (jdoe0001):

"when Joe is 12 years older" <--- meaning, Joe in 12 years from now "he will be 36" so Joe in 12 years from now, will be 36, how old is he today?

OpenStudy (jdoe0001):

as a matter of an equation, one can say that if he's "x years old" today, 12 years from now he will be "x + 12" naturally and that will EQUATE 36 since 12 years from now he'll be 36

OpenStudy (anonymous):

if joe is 36 in 12 more years, his age now can be shown as x=36-12

OpenStudy (sleepyjess):

x + 12 = 36 x = 36 - 12 x = 24